OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). (2 points) y = 2x - 7 y = - x - 2 y = 2x - 2 y = - x - 7

OpenStudy (butterflydreamer):

what is the slope of your two lines? :) looking at y = 2x + 2 ( y = mx + b, where m = slope) note: parallel lines have the same slope

OpenStudy (anonymous):

plug in 5 with X and with y in the answers they gave you

OpenStudy (anonymous):

and 3 with Y

OpenStudy (anonymous):

SO have 3 for all the Y'S and 5 for all the X's to see if they give you three

OpenStudy (anonymous):

so is either a or c correct?

OpenStudy (anonymous):

did A and C give you three ?

OpenStudy (anonymous):

3= 2 x 5 - 7 3 = - 5 - 2 3= 2 x 5 - 2 3= - 5- 7 which statement is true?

OpenStudy (butterflydreamer):

yes it will either be A or C because they both have m = 2. Now, you plug in the point (5, 3) where x = 5 and y = 3 and see which one satisfies teh equation

OpenStudy (anonymous):

is A

OpenStudy (anonymous):

so its A??

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

you seen how i got it. all you do is plug in

OpenStudy (anonymous):

but you welcome :) * best response *